OpenStudy (anonymous):

can so one help me on this i am super confused. 1. A block of material has a volume of 2 m^3. It has a mass of 1 × 10^4 kg. Calculate the density of the material sample and express your results in g/cm^3 with appropriate scientific notation.

OpenStudy (aaronq):

The difficult part in this question is the volume conversion \(1 m^3 = 1, 000 ,000 \;cm^3\) because \(1 m = 100cm,\; so\; 1^3m^3=100^3 cm^3\) kg to g is easier, 1 kg=1000g.. then once you've converted both, plug into the formula: \(density=\dfrac{mass}{volume}\)

OpenStudy (anonymous):

ok thanks

OpenStudy (aaronq):

no probs